Sally can mow a typical yard in 3 hours and joe and mo0w the same yard in 5hours. working together exactly how many hours will it take to mow 1 yard?
Answer by Alan3354(69443)

Do it like this one.
--------------------
One computer can print a company's mailing labels in 40 minutes. A second printer would take 60 minutes to print the lablels. How long would it take the two printes, operating together, to print the labels?
--------------------------------------------------------
One does 1/40 of the job per minute.
The other does 1/60 per minute.
Together, they do 1/40 + 1/60 per minute = 3/120 + 2/120 = 1/24 job per minute
--> 24 minutes per job
--------------------------
Use the shortcut, product over sum:
40*60/(40+60) = 24 minutes